OpenStudy (anonymous):

expand each logarithm: 1. log[base 8] 8 ( square root 3a^5) 2.log (s(square root 7) / t^2) 3.log [base b] 1/x Please help !

OpenStudy (anonymous):

1. \[\log_{8} 8 \sqrt{3a ^{5}}\] 2. \[\log (s \sqrt{7} \div t ^{2} )\] 3. \[\log_{b} \frac{ 1 }{ x }\]

OpenStudy (anonymous):

so i have to expand those logarithms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any thing else..??

jhonyy9 (jhonyy9):

@Kashan you dont like rewrite it here now for you can getting comments on this ,please ?

jhonyy9 (jhonyy9):

so and than the asker will can add to you medal for ... good answer

jhonyy9 (jhonyy9):

@sahura what is your opinion from this ,please ?

OpenStudy (anonymous):

@jhonyy9 what are you saying?? sorry i didn't get your point..

jhonyy9 (jhonyy9):

so if you dont like rewriting your calcules here now ?

jhonyy9 (jhonyy9):

the same how you have do it on your photo attached in form jpg.

jhonyy9 (jhonyy9):

so if we look it step by step than log(8) 8 =1 sqrt(3a^5) = (3a^5)^(1/2) @sahura do you like continue it ?

jhonyy9 (jhonyy9):

so now you need just using the property of logarithm's

OpenStudy (dumbcow):

the log properties are: \[\log (ab) = \log(a) + \log(b)\] \[\log(\frac{a}{b}) = \log(a) - \log(b)\] \[\log(a^{n}) = n*\log(a)\]

OpenStudy (anonymous):

thanks !